# The Six-Vertex Yang-Baxter Groupoid

### v2

@@ Abstract.
A parametrized Yang-Baxter equation is a map from a group to a set of R-matrices, satisfying the Yang-Baxter commutation relation. For the six-vertex model, there are two main regimes of the Yang-Baxter equation: the free-fermionic point, and everything else. For the free-fermionic point, there exists a parametrized Yang-Baxter equation with a large parameter group $\operatorname{GL}(2)\times\operatorname{GL}(1)$. For non-free-fermionic six-vertex matrices, there are also parametrized Yang-Baxter equations, but these do not account for all possible interactions. Instead we will construct a groupoid parametrized Yang-Baxter equation that does reflect all possible Yang-Baxter equations in the six-vertex model.
@@ 1. Introduction
Let $V$ be a finite-dimensional vector space. If $r,s,t\in\operatorname{End}(V\otimes V)$, the Yang-Baxter commutator is defined to be the endomorphism of $V\otimes V\otimes V$ defined by
@@
Let $G$ be a group. A (group) parametrized Yang-Baxter equation with parameter group $G$ is a map $\pi:G\longrightarrow\operatorname{End}(V\otimes V)$ such that
$\llbracket\pi(g),\pi(gh),\pi(g)\rrbracket=0.$
@@
The six-vertex model is perhaps the most basic example. Let $V=\mathbb{C}^{2}$, with standard basis $\mathbf{e}_{1},\mathbf{e}_{2}$. A six-vertex matrix is an endomorphism of $V\otimes V$, with basis $\mathbf{e}_{1}\otimes\mathbf{e}_{1},\mathbf{e}_{1}\otimes\mathbf{e}_{2}$, $\mathbf{e}_{2}\otimes\mathbf{e}_{1}$, $\mathbf{e}_{2}\otimes\mathbf{e}_{2}$, having a matrix of the form
@@
We will always consider only $u$ with $c_{1}(u),c_{2}(u)$ nonzero. The entries in the matrix may be referred to as Boltzmann weights due to the origin of the subject in statistical mechanics [1].
Let us give two examples of parametrized Yang-Baxter equations.
First, let $q_{1}$ and $q_{2}$ be fixed, nonzero complex numbers. For $z_{1},z_{2},w\in\mathbb{C}^{\times}$, define a six-vertex matrix $R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ to be the matrix with the following Boltzmann weights:
@@
$\llbracket R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w_{1}),R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1}z_{3},z_{2}z_{4},w_{1}w_{2}),R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{3},z_{4},w_{2})\rrbracket=0.$
@@
A second example is, with $z_{1},z_{2},w\in\mathbb{C}^{\times}$, the $R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ to be the matrix with the following Boltzmann weights:
@@
$\llbracket R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{1},z_{2},w_{1}),R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{1}z_{3},z_{2}z_{4},w_{1}w_{2}),R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{3},z_{4},w_{2})\rrbracket=0.$
@@
These two examples are very similar: the weights are the same except for that $a_{2}$ entries, which differ. The weights $R^{\operatorname{ff}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ are free-fermionic, meaning that they satisfy the condition
@@
On the other hand the weights $R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ are not free fermionic, but instead satisfy the condition $a_{1}(R)=a_{2}(R)$.
@@
Despite the similarity between the above two examples, there are nevertheless important differences between the free-fermionic regime and the non-free-fermionic regime in the six-vertex model. The free-fermionic example $R^{\operatorname{ff}}_{q_{1},q_{2}}$ can be extended to a parametrized Yang-Baxter equation with a larger group $\operatorname{GL}(2,\mathbb{C})\times\operatorname{GL}(1,\mathbb{C})$. In contrast $R^{\operatorname{cf}}_{q_{1},q_{2}}$ does not extend to a larger group.
@@
As was shown by Drinfeld [5] and Jimbo [7], solutions to the Yang-Baxter equation are explained by the theory of quantum groups. For non-free-fermionic parametrized Yang-Baxter equations such as $R^{\operatorname{cf}}$, the relevant quantum group is $U_{q}(\mathfrak{g}\mathfrak{l}_{2})$ or its affinization, or Drinfeld twists, including two-parameter quantum groups. For free-fermionic Yang-Baxter equation such as $R^{\operatorname{ff}}$, the relevant quantum group is the superalgebra $U_{q}(\mathfrak{s}\mathfrak{l}(1|1))$, or its affinization ([2, 3]).
@@
Now there are substantial differences between the representation theories of $U_{q}(\mathfrak{g}\mathfrak{l}_{2})$ and $U_{q}(\mathfrak{s}\mathfrak{l}(1|1))$. For example $U_{q}(\mathfrak{s}\mathfrak{l}(1|1))$ has many irreducible two-dimensional representations, accounting for the large $\operatorname{GL}(2)\times\operatorname{GL}(1)$-parametrized Yang-Baxter equation in the free-fermionic case. In contrast, this is not true for $U_{q}(\mathfrak{sl}_{2})$, which has only one. (Its affinization has a one-parameter family of two-dimensional representations.) Correspondingly, we do not find such a large (group) parametrized Yang-Baxter equation in the non-free-fermionic case.
@@
Moreover, the group parametrized Yang-Baxter equations in the non-free-fermionic case do not account for all possible interactions between six-vertex matrices. To demonstrate this, let us consider the parametrized Yang-Baxter equation $R^{\operatorname{cf}}_{q_{1},q_{2}}$. We choose an element $r=R^{\operatorname{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ of this group. There are always six-vertex matrices $u$ and $v$ that are outside the group such that $\llbracket u,v,r\rrbracket=0$. Indeed, by Corollary 3.2 below, if $u$ is any six-vertex matrix such that $\Delta(u)=\Delta(r^{\ast})=(\frac{q_{1}+q_{2}}{q_{2}},\frac{q_{1}+q_{2}}{q_{1}})$, then there exists a $u$ such that $\llbracket u,v,r\rrbracket=0$. The condition on $u$ restricts it to a four-dimensional variety, but the group of matrices $R^{\operatorname{cf}}_{q_{1},q_{2}}$ is three-dimensional, so $u$ and $v$ that are outside the group certainly exist.
However if we expand the concept of a parametrized Yang-Baxter equation to allow the parameter object to be a groupoid, rather than a group, then we will show that all such interactions can be accounted for. Recall that a groupoid $\mathfrak{G}$ is a set with an associative composition law, which we will denote $\star$, that is only partially defined. Thus if $u,v\in\mathfrak{G}$, it may or may not be the case that $u,v$ is defined.
@@
Thus we may define a groupoid parametrized Yang-Baxter equation to be a map $\pi:\mathfrak{G}\to\operatorname{End}(V\otimes V)$, where $\mathfrak{G}$ is a groupoid and $V$ a vector space, to be a map such that if $g,h\in\mathfrak{G}$ are such that $g\star h$ is defined, then $\llbracket\pi(g),\pi(g\star h),\pi(h)\rrbracket=0$. We will construct such a groupoid parametrized Yang-Baxter equation. The groupoid $\mathfrak{G}$ will decompose into a disjoint union of a free-fermionic part $\mathfrak{G}_{\operatorname{ff}}$, which is actually the group $\operatorname{GL}(2)\times\operatorname{GL}(1)$, and a more interesting non-free-fermionic part $\mathfrak{G}_{\operatorname{nf}}$, which is a groupoid.
@@
Turning to the non-free-fermionic groupoid, with $V=\mathbb{C}^{2}$, the map $\pi:\mathfrak{G}_{\operatorname{nf}}\to\operatorname{End}(V\otimes V)$ is not an isomorphism onto its image. The set $\pi(\mathfrak{G}_{\operatorname{nf}})$ is a specific set $\overline{\Omega}$ of matrices in $\operatorname{End}(V\otimes V)$. The set $\overline{\Omega}$ can be thought of as a quasi-affine variety, and it has a Zariski topology. The set $\pi:\mathfrak{G}_{\operatorname{nf}}\rightarrow\overline{\Omega}$ is a morphism that is a birational equivalence. Essentially, $\mathfrak{G}_{\operatorname{nf}}$ is obtained by “blowing up” certain lower-dimensional boundary components in a specific way. We refer to the last section for details.
The set $\overline{\Omega}$ has an open subset $\Omega^{\circ}$, and three boundary components. A special role is played by two functions on $\overline{\Omega}$. These are:
$\Delta_{1}(u)=\frac{a_{1}(u)a_{2}(u)+b_{1}(u)b_{2}(u)-c_{1}(u)c_{2}(u)}{a_{1}(u)b_{1}(u)},$
$\Delta_{2}(u)=\frac{a_{1}(u)a_{2}(u)+b_{1}(u)b_{2}(u)-c_{1}(u)c_{2}(u)}{a_{2}(u)b_{2}(u)}.$
Let $\Delta(u)=(\Delta_{1}(u),\Delta_{2}(u))$. This is a function on $\Omega^{\circ}$ that is not defined on certain boundary components, but it can be defined on the blowup $\mathfrak{G}_{\operatorname{nf}}$. Now if $u,v\in\mathfrak{G}_{\operatorname{nf}}$, then $u\star v$ is defined if and only if $\Delta(u)=\Delta(v^{\prime})$, and if $w=u\star v$ then $\Delta(w)=\Delta(v)$.
@@
Define $\Delta_{0}(u)=\Delta_{1}(u)\Delta_{2}(u)$. If $\delta_{0}$ is given, let
$\mathfrak{G}_{\operatorname{nf}}(d_{0})=\{u\in\mathfrak{G}_{\operatorname{nf}}|\Delta_{0}(u)=\delta_{0}\}.$
Then we have a further decomposition
$\mathfrak{G}_{\operatorname{nf}}=\bigsqcup_{d_{0}}\mathfrak{G}_{\operatorname{nf}}(d_{0})$
Indeed, Proposition 3.4 shows that if $\llbracket u,w,v\rrbracket=0$ then $\Delta_{0}(u)=\Delta_{0}(w)=\Delta_{0}(v)$. This further decomposition of the groupoid is certainly important but has no role in the proofs.
One may speculate that there are other Yang-Baxter equations parametrized by groupoids. Another paper relating parametrized Yang-Baxter equations to groupoids is Felder and Ren [6]. However their use of a groupoid is different from ours.
@@
We thank Amol Aggarwal, Andrew Hardt and Travis Scrimshaw for helpful conversations about the groupoid.

### v3

@@ Abstract.
A parametrized Yang-Baxter equation is usually defined to be a map from a group to a set of R-matrices, satisfying the Yang-Baxter commutation relation. These are a mainstay of solvable lattice models. We will show how the parameter space can sometimes be enlarged to a groupoid, and give two examples of such groupoid parametrized Yang-Baxter equations, within the six vertex model. A groupoid parametrized Yang-Baxter equation consists of a groupoid $\mathfrak{G}$ together with a map $\pi:\mathfrak{G}\to\operatorname{End}(V\otimes V)$ for some vector space $V$ such that the Yang-Baxter commutator $\llbracket\pi(u),\pi(w),\pi(v)\rrbracket=0$ if $u,v\in\mathfrak{G}$ are such that the groupoid composition $w=u\star v$ is defined. An important role is played by an object map $\Delta:\mathfrak{G}\to M$ for some set $M$ such that $\Delta(u)=\Delta(v^{\prime})$, $\Delta(w)=\Delta(v)$ and $\Delta(w^{\prime})=\Delta(u^{\prime})$, where $v\mapsto v^{\prime}$ is the groupoid inverse map.
There are two main regimes of the six-vertex model: the free-fermionic point, and everything else. For the free-fermionic point, there exists a parametrized Yang-Baxter equation with a large parameter group $\operatorname{GL}(2)\times\operatorname{GL}(1)$. For non-free-fermionic six-vertex matrices, there are also well-known (group) parametrized Yang-Baxter equations, but these do not account for all possible interactions. Instead we will construct a groupoid parametrized Yang-Baxter equation that accounts for essentially all possible Yang-Baxter equations in the six-vertex model. We will also exhibit a separate groupoid for the five-vertex model. We will show how to construct solvable lattice models based on groupoid parametrized Yang-Baxter equations.
@@ 1. Introduction
Let $V$ be a finite-dimensional vector space. If $r,s,t\in\mathrm{End}(V\otimes V)$, the Yang-Baxter commutator is defined to be the endomorphism of $V\otimes V\otimes V$ defined by
@@
Let $G$ be a group. A (group) parametrized Yang-Baxter equation with parameter group $G$ is a map $\pi:G\longrightarrow\mathrm{End}(V\otimes V)$ such that
(1) $\llbracket\pi(g),\pi(gh),\pi(g)\rrbracket=0.$
@@
Group parametrized Yang-Baxter equations are a mainstay of solvable lattice model theory, underlying much work starting with Baxter’s treatment of the field free six- and eight-vertex model [1]. In this paper we propose a generalization, replacing the group $G$ by a groupoid. We review groupoids in Section 2.
To define the notion of a groupoid parametrized Yang-Baxter equation, let $\mathfrak{G}$ be a groupoid, and $\pi:\mathfrak{G}\longrightarrow\operatorname{End}(V\otimes V)$ a map such that if $u,v\in G$ and if the product $u\star v$ is defined in $\mathfrak{G}$, then $\llbracket\pi(u),\pi(u\star v),\pi(v)\rrbracket=0$. We will further assume that there exists a set $M$ and a map $\Delta:\mathfrak{G}\longrightarrow M$ such that $u\star v$ is defined if and only if $\Delta(u)=\Delta(v^{\prime})$, where $v^{\prime}$ is the groupoid inverse of $v$, and if $w=u\star v$ then
(2) $\Delta(u)=\Delta(v^{\prime}),\qquad\Delta(w)=\Delta(v),\qquad\Delta(w^{\prime})=\Delta(u^{\prime}).$
If $\mathfrak{G}$ is a group, we may take $M$ to be a single point, so the map $\Delta$ contains no information. We will call the function $\Delta:\mathfrak{G}\longrightarrow M$ the object map.
Remark 1.
The existence of the map $\Delta$ follows from the first part of the definition. Indeed, by Proposition 2.5 every groupoid is associated with a groupoid category $\mathcal{C}$. Then $M$ can be taken to be the set of objects of $\mathcal{C}$, while the groupoid $\mathfrak{G}$ is the set of all morphisms in $\mathcal{C}$. The map $\Delta:\mathfrak{G}\longrightarrow M$ then associates with a morphism $v$ its source object $\Delta(v)$. The target of the morphism is $\Delta(v^{\prime})$. The groupoid composition law is the composition in the category $\mathcal{C}$, and all three identities in (2) then have interpretations. The first identity says that if $u$ and $v$ are morphisms the composition $u\circ v$ is defined if and only if the source of $u$ is the target of $v$, and the other identities say that the source of $u\circ v$ is the source of $v$, and that the target of $u\circ v$ is the target of $u$. Although the existence of the map $\Delta$ is thus automatic, in practice $M$ is a concrete set and the map $\Delta$ plays an important role in the construction of the groupoid. Therefore we are making it a part of the definition.
In the simplest case the groupoid is just a disjoint union of groups. In this case, $M$ could be any set that is in bijection with the set of groups in the disjoint union. If this is so, then $\Delta(v)=\Delta(v^{\prime})$, and $\llbracket u,w,v\rrbracket=0$ implies that $\Delta(u)=\Delta(v)=\Delta(w)$.
In this paper we will give two groupoid parametrized Yang-Baxter equations, related to the (non-free-fermionic) six-vertex model, and the five-vertex models. These groupoids are not disjoint unions of groups.
Both of these groupoids contain well-known subgroups that are the basis of interesting Yang-Baxter equations with applications to solvable lattice models. So to demonstrate that groupoid parametrized are potentially important, in Section 8 we will consider some solvable lattice models based on an arbitrary groupoid parametrized Yang-Baxter equations. We will see that such lattice models can be constructed with both row and column solvability. These are similar to well-known models (for example [1, 10, 4]) which indeed correspond to subgroups of the groupoid. But if we work with a groupoid parametrized Yang-Baxter equation in which the groupoid is not a disjoint union of groups, the examples in Section 8 are more general.
1.1. The Six-Vertex Model
Let $V=\mathbb{C}^{2}$, with standard basis $\mathbf{e}_{1},\mathbf{e}_{2}$. A six-vertex matrix is an endomorphism of $V\otimes V$, with basis $\mathbf{e}_{1}\otimes\mathbf{e}_{1},\mathbf{e}_{1}\otimes\mathbf{e}_{2}$, $\mathbf{e}_{2}\otimes\mathbf{e}_{1}$, $\mathbf{e}_{2}\otimes\mathbf{e}_{2}$, having a matrix of the form
@@ 1.1. The Six-Vertex Model
Furthermore we require that $c_{1}(u),c_{2}(u)$ is nonzero. The entries in the matrix maeq:deltatry be referred to as Boltzmann weights due to the origin of the subject in statistical mechanics [1].
Let us give two examples of group parametrized Yang-Baxter equations in the six-vertex model. First, let $q_{1}$ and $q_{2}$ be fixed, nonzero complex numbers. For $z_{1},z_{2},w\in\mathbb{C}^{\times}$, define a six-vertex matrix $R^{\mathrm{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ to be the matrix with the following Boltzmann weights:
@@
$\llbracket R^{\mathrm{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w_{1}),R^{\mathrm{cf}}_{q_{1},q_{2}}(z_{1}z_{3},z_{2}z_{4},w_{1}w_{2}),R^{\mathrm{cf}}_{q_{1},q_{2}}(z_{3},z_{4},w_{2})\rrbracket=0.$
@@
A second example is, with $z_{1},z_{2},w\in\mathbb{C}^{\times}$, the $R^{\mathrm{ff}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ to be the matrix with the following Boltzmann weights:
@@
$\llbracket R^{\mathrm{ff}}_{q_{1},q_{2}}(z_{1},z_{2},w_{1}),R^{\mathrm{ff}}_{q_{1},q_{2}}(z_{1}z_{3},z_{2}z_{4},w_{1}w_{2}),R^{\mathrm{ff}}_{q_{1},q_{2}}(z_{3},z_{4},w_{2})\rrbracket=0.$
@@
These two examples are very similar: the weights are the same except for that $a_{2}$ entries, which differ. The weights $R^{\mathrm{ff}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ are free-fermionic, meaning that they satisfy the condition
@@
On the other hand the weights $R^{\mathrm{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ are not free fermionic, but instead satisfy the condition $a_{1}(R)=a_{2}(R)$.
@@
Despite the similarity between the above two examples, there are nevertheless important differences between the free-fermionic regime and the non-free-fermionic regime in the six-vertex model. The free-fermionic example $R^{\mathrm{ff}}_{q_{1},q_{2}}$ can be extended to a parametrized Yang-Baxter equation with a larger group $\mathrm{GL}(2,\mathbb{C})\times\mathrm{GL}(1,\mathbb{C})$. In contrast $R^{\mathrm{cf}}_{q_{1},q_{2}}$ does not extend to a properly larger connected group.
@@
The dichotomy between the free-fermionic and non-free-fermionic cases can be explained as follows. Drinfeld [6] and Jimbo [8] showed that solutions to the Yang-Baxter equation come from quantum groups. For non-free-fermionic parametrized Yang-Baxter equations such as $R^{\mathrm{cf}}$, the relevant quantum group is $U_{q}(\mathfrak{g}\mathfrak{l}_{2})$ or its affinization, or Drinfeld twists, including two-parameter quantum groups. For free-fermionic Yang-Baxter equation such as $R^{\mathrm{ff}}$, the relevant quantum group is the superalgebra $U_{q}(\mathfrak{s}\mathfrak{l}(1|1))$, or its affinization ([2, 3, 5]).
@@
Differences between the representation theories of $U_{q}(\mathfrak{gl}_{2})$ and $U_{q}(\mathfrak{sl}(1|1))$ explain some of the differences between these cases. Thus $U_{q}(\mathfrak{sl}(1|1))$ has many irreducible two-dimensional representations, accounting for the large $\mathrm{GL}(2)\times\mathrm{GL}(1)$-parametrized Yang-Baxter equation in the free-fermionic case. In contrast, this is not true for $U_{q}(\mathfrak{sl}_{2})$, which has essentially only one. (Its affinization has a one-parameter family of two-dimensional representations.)
@@
The group parametrized Yang-Baxter equations in the non-free-fermionic case do not account for all possible interactions between six-vertex matrices. To demonstrate this, let us consider the parametrized Yang-Baxter equation $R^{\mathrm{cf}}_{q_{1},q_{2}}$. We choose an element $r=R^{\mathrm{cf}}_{q_{1},q_{2}}(z_{1},z_{2},w)$ of this group. There are always six-vertex matrices $u$ and $v$ that are outside the group such that $\llbracket u,v,r\rrbracket=0$. Indeed, by Corollary 3.2 below, if $u$ is any six-vertex matrix such that $\Delta(u)=\Delta(r^{\ast})=(\frac{q_{1}+q_{2}}{q_{2}},\frac{q_{1}+q_{2}}{q_{1}})$, then there exists a $u$ such that $\llbracket u,v,r\rrbracket=0$. The condition on $u$ restricts it to a four-dimensional variety, but the group of matrices $R^{\mathrm{cf}}_{q_{1},q_{2}}$ is three-dimensional, so $u$ and $v$ that are outside the group certainly exist.
Theorem.
There is a groupoid parametrized Yang-Baxter equation $\pi:\mathfrak{G}\to\mathrm{End}(V\otimes V)$ with $V=\mathbb{C}^{2}$ with object map $\Delta:\mathfrak{G}\longrightarrow(\mathbb{C}^{\times})^{2}$ such that if $u$ and $v$ are six-vertex matrices in general position then there exists a six-vertex matrix $w$ such that $\llbracket u,w,v\rrbracket=0$ if and only if there exist $g,h\in\mathfrak{G}$ such that $u=\pi(g)$ and $v=\pi(h)$, and $g\star h$ is defined. Then $w=\pi(g\star h)$ is the solution.
@@
Thus this groupoid parametrized Yang-Baxter equation accounts for all Yang-Baxter identities $\llbracket r,s,t\rrbracket=0$ with $r,s,t$ in an open subset $S^{\circ}$ of six-vertex matrices. A disjoint union of groupoids is a groupoid. The groupoid $\mathfrak{G}$ decomposes into the disjoint union of a free-fermionic part $\mathfrak{G}_{\mathrm{ff}}$, which is actually the group $\mathrm{GL}(2)\times\mathrm{GL}(1)$, and a more interesting non-free-fermionic part $\mathfrak{G}_{\mathrm{nf}}$, which is a groupoid, but not a disjoint union of groups.
Remark 2.
For $r,s,t$ outside $S^{\circ}$, there may be further boundary groupoids that we do not fully analyze. Particularly, there are five-vertex matrices which have themeselves important applications. We will prove in Section 7 that the matrices with $b_{2}=0$ can be organized into an interesting groupoid. The matrices with $b_{1}=0$ similarly fit into a groupoid. These lower-dimensional groupoids are not included in $\mathfrak{G}$.
@@
In this introductory section we will describe the groupoid $\mathfrak{G}$ in rough terms, as an outline of the somewhat technical construction that will follow in later sections. The map $\pi:\mathfrak{G}_{\mathrm{nf}}\to\mathrm{End}(V\otimes V)$ is not an isomorphism onto its image. The set $\pi(\mathfrak{G}_{\mathrm{nf}})$ is a specific set $\bar{\Omega}$ of matrices in $\mathrm{End}(V\otimes V)$. The set $\bar{\Omega}$ can be thought of as a quasi-affine variety, and it has a Zariski topology. The set $\pi:\mathfrak{G}_{\mathrm{nf}}\rightarrow\bar{\Omega}$ is a morphism that is a birational equivalence. Essentially, $\mathfrak{G}_{\mathrm{nf}}$ is obtained by “blowing up” certain lower-dimensional boundary components.
The object map $\Delta$ takes values in $M=(\mathbb{C}^{\times})^{2}$. If $u\in\Omega$ define
$N(u)=a_{1}(u)a_{2}(u)+b_{1}(u)b_{2}(u)-c_{1}(u)c_{2}(u)$
$\Delta_{1}(u)=\frac{N(u)}{a_{1}(u)b_{1}(u)},\qquad\Delta_{2}(u)=\frac{N(u)}{a_{2}(u)b_{2}(u)}.$
We find that
$\Delta_{1}(u^{\prime})=\frac{N(u)}{a_{1}(u)b_{1}(u)},\qquad\Delta_{2}(u^{\prime})=\frac{N(u)}{a_{2}(u)b_{2}(u)}.$
Define $\Delta(u)=(\Delta_{1}(u),\Delta_{2}(u))$. This function is holomorphic on $\Omega^{\circ}$ but it cannot be extended to all of $\overline{\Omega}$ since the numerator and denominator can both vanish. This is mitigated by blowing up two boundary components, and $\Delta$ extends to a continuous function on all of $\mathfrak{G}$.
Although the six-vertex groupoid is not a disjoint union of groups, it can still be decomposed into a disjoint union of smaller groupoids that we will call blocks.
@@
Define $\Delta_{0}(u)=\Delta_{1}(u)\Delta_{2}(u)$. It may be checked that $\Delta_{0}(u)=\Delta_{0}(u^{\prime})$. If $\delta_{0}\in\mathbb{C}^{\times}$ is given, define the block
$\mathfrak{G}_{\mathrm{nf}}(d_{0})=\{u\in\mathfrak{G}_{\mathrm{nf}}|\Delta_{0}(u)=\delta_{0}\}.$
Note that if $u$ and $v$ lie in different blocks then $\Delta(u)=\Delta(v^{\prime})$ is impossible since this would imply that $\Delta_{0}(u)=\Delta_{0}(v^{\prime})$. Thus if $u,v,w\in\mathfrak{G}$ and $\llbracket u,w,v\rrbracket=0$ all three $u,v,w$ must lie in the same block. Thus we have a further decomposition
$\mathfrak{G}_{\mathrm{nf}}=\bigsqcup_{d_{0}}\mathfrak{G}_{\mathrm{nf}}(d_{0}).$
Another paper relating parametrized Yang-Baxter equations to groupoids is Felder and Ren [7]. However their use of a groupoid is different from ours.
@@
We thank Amol Aggarwal, Ben Brubaker, Valentin Buciumas, Henrik Gustafsson, Andrew Hardt and Travis Scrimshaw for helpful conversations and comments.